OpenStudy (rose4825):

Please check: What is the first step in solving x^2 + 2x + 3 = 0 by completing the square? A. Subtract 3 from each side of the equation. B. Add 3 to each side of the equation. C. Factor the left side. D. Divide both sides of the equation by 2. My answer: A

OpenStudy (rose4825):

@wolf1728

OpenStudy (wolf1728):

I'd say you are right it is A

OpenStudy (rose4825):

Thank you so much

OpenStudy (wolf1728):

glad to help :-)

OpenStudy (3mar):

@rose4825 @wolf1728 what after "Subtract 3 from each side of the equation"?

OpenStudy (wolf1728):

The next step is to divide the equation by the coefficient of x^2 and since it is 1 that doesn't have to be done.

OpenStudy (rose4825):

It only asks for the first step @3mar.

OpenStudy (wolf1728):

If you need to know all the steps of completing the square, try this web page http://www.1728.org/quadr2.htm read the first half.

OpenStudy (3mar):

So you mean that: \[x^2 + 2x + 3 = 0\] \[x^2 + 2x + 3 -3= 0-3\] \[x^2 + 2x = -3\] \[\frac{1 }{ 1 } x^2 + \frac{ 2 }{ 1 } x = \frac{ -3 }{ 1 }\] \[x^2 + 2x = -3\] then add and subtract \((\frac{ 2 }{ 2 })^2\) to the left-hand side \[x^2 + 2x +1-1= -3\] \[(x^2 + 2x +1)-1= -3\] \[(x+1)^2-1= -3\] So that is the aim of us,,,right?

OpenStudy (wolf1728):

x^2 + 2x = -3 x^2 + 2x +1= -3 +1 (x+1)^2 = -2 x + 1 = sqrt(-2) x = sqrt(-2) -1

OpenStudy (3mar):

But I think: \[x^2 + 2x +1= -3 +1\\(x+1)^2=-2\\ \sqrt{(x+1)^2}=\sqrt{-2}\\x+1=\pm \sqrt{2}~i\\x=-1\pm \sqrt{2}~i\] therefore: \[\Huge {x_1=-1+ \sqrt{2}~i\\or\\x_2=-1- \sqrt{2}~i}\] that is just in case he wants to solve this equation, not to complete the square!! agree with me @wolf1728

OpenStudy (3mar):

Let's involve @rose4825 also!

OpenStudy (wolf1728):

i agree with your solution

OpenStudy (wolf1728):

Hmm it seems rose is offline

OpenStudy (3mar):

That is good,,,and don't forget that the output of the square root results in two quantities (negative and positive).... That point always makes a confusion in results!

OpenStudy (wolf1728):

So, I guess you know about completing the square pretty well.

OpenStudy (3mar):

yes. Praise for Allah! and all of us learn and add new experience every day...that what makes us grow..

OpenStudy (wolf1728):

I guess so. This place (openstudy) is good for keeping the mind sharp.

OpenStudy (3mar):

I agree with you...but firstly it needs to be controlled and moderated by some serious and honest stuff...

OpenStudy (wolf1728):

I always thought that this place was fairly well-moderated

OpenStudy (rose4825):

Thank you both. I did log off because I had to go to bed, but I understand what you guys mean.

OpenStudy (3mar):

I am happy to hear that. What matters is that you got the idea!